Find an equation of the hyperbola that satisfies the given conditions.
Vertices (−1, 7) and (−1, 3), foci (−1, 9) and (−1, 1)

Answer by Boreal(15235)

The center is midway between the vertices and is (-1,5), It is also midway between both foci.
c=4, distance of each focus from the center
a=2 distance of each vertex from the center.
c^2=a^2+b^2
b^2=12. Don't need b itself, since the equation has b^2 only.
y is added, since the vertices are above one another.
{(y-5)^2/16} - {(x+1)^2/12} =1